jonathan14:

A cylindrical fish tank at the zoo has a capacity of 8139 cubic feet.What is the height of the tank if its radius is 18 feet? Use and round to the nearest foot. pie is 3.14

Florisalreadytaken:

A cylindrical fish tank at the zoo has a capacity of 8139 cubic feet.What is the height of the tank if its radius is 18 feet? Use and round to the nearest foot. pie is 3.14 Well We know that \( \boxed{V_{tank}= 8139 \ ft^2 }\) & \(\boxed{ r=18 \ ft }\) where \( V_{tank}= \pi r^2 \times h \) we need to find the height \((h)\) so solving for it: \( \boxed{ \dfrac{V_{tank}}{\pi r^2}= h} \) Just put the values in at this point: \( \dfrac{ 8139}{3.14 \times 18^2}= h \) Just use a calculator to find the value.

Florisalreadytaken:

Also, why did the person who created this question name it "Pie" haha !